undeadknight26 (undeadknight26):

Compare the functions shown below: f(x) = 3 cos 2x + 4 g(x) cosine graph with points at 0, 3 and pi over 2, 0 and pi, negative 3 and 3 pi over 2, 0 and 2 pi, 3 h(x) x y -6 -11 -5 -6 -4 -3 -3 -2 -2 -3 -1 -6 0 -11

undeadknight26 (undeadknight26):

@sleepyjess @shifuyanli @dan815 @Freez @geerky42 @gabbyalicorn @horsegirl325 @JoannaBlackwelder @wwhitlock @texaschic101 @campbell_st @calculusxy

OpenStudy (gabbyalicorn):

Sorry math is not my best subject

geerky42 (geerky42):

What do you think? @undeadknight26

geerky42 (geerky42):

Maximum possible value of \(\cos\theta\) is 1, right?

OpenStudy (joannablackwelder):

Can you tell me what the max values of each function are?

undeadknight26 (undeadknight26):

Oh mai i forgot to say what i think /.\ Mighty embarrassing.... I thought it was f(x)! Since its maximum y value was (0,7)

geerky42 (geerky42):

Yes maximum of f(x) is 7

OpenStudy (wwhitlock):

Looks like math to me.

geerky42 (geerky42):

It is math lol...

geerky42 (geerky42):

Can you tell me the maximum of g(x) and h(x)?

OpenStudy (joannablackwelder):

Yep, max value of f(x) is 7 :-)

undeadknight26 (undeadknight26):

For me f(x) = 7 g(x) = 3 h(x) = -2. I just want to know if i am right with this...

geerky42 (geerky42):

You are correct.

geerky42 (geerky42):

Therefore answer is f(x)

undeadknight26 (undeadknight26):

Yosh thanks guys! Care to check like 4 more?

geerky42 (geerky42):

Sure

undeadknight26 (undeadknight26):

I think this one is f(x) again

geerky42 (geerky42):

Yep

undeadknight26 (undeadknight26):

f(x) = 5 g(x) = 2 h(x) = 1

geerky42 (geerky42):

Actually h(x) is -2

geerky42 (geerky42):

y-intercept is when x=0, right? so just plug in x=0 then solve for y.

geerky42 (geerky42):

\[3\sin(2(0) + \pi )-2 = 3\sin(\pi)-2 = 3(0)-2 = \boxed{-2}\]

undeadknight26 (undeadknight26):

Oh yosh that makes sense :O

undeadknight26 (undeadknight26):

I think its D

undeadknight26 (undeadknight26):

I think this one is C

undeadknight26 (undeadknight26):

I think this one is A

geerky42 (geerky42):

All correct.